OpenStudy (anonymous):

Factor: x^2-8x+16 I need to factor this, but I am not understanding what to do. Please use clear instructions. Here are some of the possible answers: (x+8)^2 (x-8)^2 (x+4)^2 (x-4)^2 I think it's the last one but again, I have no idea how to do this. Help?

OpenStudy (jango_in_dtown):

x^2-8x+16 imply x^2-4x-4x+16=0 imply x(x-4)-4(x-4)=0 imply (x-4)(x-4)=0 imply (x-4)^2=0

OpenStudy (anonymous):

Okay, @jango_IN_DTOWN I kind of get what you're saying, can you kind of explain whats happening in each step?

OpenStudy (jango_in_dtown):

see actually see it is of the form (a-b)^2

OpenStudy (jango_in_dtown):

(a-b)^2=a^2-2ab+b^2

OpenStudy (anonymous):

Okay, I get the first part, thanks! @jango_IN_DTOWN Can you explain what you did with the 16?

OpenStudy (jango_in_dtown):

yeah i wrote 16=4*4 and took one 4 common from it and the previous term

OpenStudy (anonymous):

mmh! Okay that makes sense! Now what about the second to last step? How did you get rid of the x and the -4 to get (x-4)(x-4) ? @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

see ax+bx=(a+b)x I used this property

OpenStudy (anonymous):

Okay, last question, do you know what property it's called? @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

distributive property you may check let a=2, b=1,c=3 then a(b+c)=2(1+3)=2*4=8 and ab+ac=2*1+2*3=2+6=8

OpenStudy (anonymous):

Thanks so much, you're awesome! @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

n.p.;)